Foreign Exchange. Determine the size of the government purchase.
D: P = 680 - 2Q
S: P = 40 + 2Q
Target: P = 500
1
Expert's answer
2016-05-25T12:42:03-0400
D: P = 680 - 2Q
S: P = 40 + 2Q
In equilibrium Ps = Pd and Qs = Qd, so:
680 - 2Q = 40 + 2Q,
4Q = 640,
Q = 160 units.
P = 680 - 2*160 = $360.
If the target price is P = 500, then the quantity demanded is:
500 = 680 - 2Q,
2Q = 180,
Q = 90 units.
The quantity supplied is:
500 = 40 + 2Q,
2Q = 460,
Q = 230 units.
So, there is a surplus of 230 - 90 = 140 units and the size of the government purchase is G = 500*140 = $70,000.